Johnson Company purchased a computer system for $66,800 on January 1, 2019. It was depreciated based on a 6-year life and an $17,000 salvage value. On January 1, 2021, Johnson revised these estimates to a total useful life of 4 years and a salvage value of $10,000. Johnson uses straight-line depreciation.

Prepare Johnson's entry to record 2021 depreciation expense
Answer #1

Depreciation Expense per year = (66,800-17,000)/6 = 8300 per year

Book value on Jan 01,2021

= 66,800 - (8300*2) = 50,200

New depreciation expense = (50,200-10,000)/(4-2) = 20,100

Depreciation expense 20100
Accumulated Depreciation - computer 20100